# ReconX: Reconstruct Any Scene from Sparse Views with Video Diffusion Model

### v2

@@ Appendix D More Results and Analysis
To further investigate the generative capabilities of our framework and demonstrate its extrapolation potential, we conduct experiments by conditioning on the first and an intermediate frame of the target video with a new tuning version of the video diffusion model in ReconX by only moving the last frame to the intermediate position. In this setup, frames between the first and intermediate images correspond to view interpolation, while frames beyond the intermediate image correspond to extrapolation.
@@
Such extrapolation ability allows us to even recover a 360-degree scene from only two sparse views. Specifically, we adopt an incremental generation approach shown in Figure 9. Given two initial input images (i.e., input image 1 and input image 2 in Figure 9), we first generate a video sequence divided into four segments: input image 1, view interpolation, input image 2, and view extrapolation. From this generated frame sequence, we select two images—one from the interpolation part and another one from the extrapolation part. These two images function as a sliding window, and repeat the generation process, progressively advancing with each iteration. This approach allows our framework to autoregressively generate a much longer 360-degree panoramic sequence while maintaining a limited-length video frame window. In the final iteration, we select one image from the video generated in the previous step (i.e., select image 1 in Figure 9) and pair it with the original first input image (i.e., input image 1 in Figure 9) as the input pair. This ensures a seamless connection back to the starting view, completing a full 360-degree scene reconstruction shown in Figure 11. This incremental approach demonstrates the strong generative extrapolation potential of our method.
Figure 12: Qualitative results of our ReconX on outdoor scenes Ling et al. (2024).

### v3

@@ V-D Evaluation of Extrapolation Ability
As the position of our conditional images in ReconX is inherently flexible, allowing us to unleash more extrapolation capability by adjusting the placement of the conditional images. To further investigate the generative capabilities of our framework and demonstrate its extrapolation potential, we conduct experiments by conditioning on the first and an intermediate frame of the target video with a new tuning version of the video diffusion model in ReconX by only moving the last frame to the intermediate position. In this setup, frames between the first and intermediate images correspond to view interpolation, while frames beyond the intermediate image correspond to extrapolation.
@@
Such extrapolation ability allows us to even recover a 360-degree scene from only two sparse views. Specifically, we adopt an incremental generation approach shown in Fig. 8. Given two initial input images (i.e., input image 1 and input image 2 in Fig. 8), we first generate a video sequence divided into four segments: input image 1, view interpolation, input image 2, and view extrapolation. From this generated frame sequence, we select two images—one from the interpolation part and another one from the extrapolation part. These two images function as a sliding window, and repeat the generation process, progressively advancing with each iteration. This approach allows our framework to autoregressively generate a much longer 360-degree panoramic sequence while maintaining a limited-length video frame window. In the final iteration, we select one image from the video generated in the previous step (i.e., select image 1 in Fig. 8) and pair it with the original first input image (i.e., input image 1 in Fig. 8) as the input pair. This ensures a seamless connection back to the starting view, completing a full 360-degree scene reconstruction shown in Fig. 9. This incremental approach demonstrates the strong generative extrapolation potential of our method.
V-E Ablation Study and Analysis
We carry out ablation studies on RealEstate10K to analyze the design of our ReconX framework in Fig. 10 and extends the ablation to contain all different combinations of leaving out individual components in Table V. A naive combination of pretrained video diffusion model and Gaussian Splatting is regarded as the “base”. Specifically, we ablate on the following aspects of our method: 3D structure condition, DUSt3R initialization, confidence-aware optimization, and LPIPS loss. The results indicate that the omission of any of these elements leads to a degradation in terms of quality and consistency. Notably, the basic combination of original video diffusion model and 3DGS leads to significant distortion of the scene. The absence of 3D structure condition causes inconsistent generated frames especially in distant input views, resulting in blur and artifact issues. The lack of confidence-aware optimization leads to suboptimal results in some local detail areas. Adding LPIPS loss in confidence-aware optimization would provide clearer rendering views. This illustrates the effectiveness of our overall framework, which drives generalizable and high-fidelity 3D reconstruction given only sparse views as input.
Moreover, we ablate the impact of DUSt3R and video diffusion priors in Fig. 11. Although the point cloud may not include enough high-quality information, such coarse 3D structure is sufficient to guide the video diffusion in our ReconX to fill in the distortions, occlusions, or missing regions. This demonstrates that our ReconX has learned a comprehensive understanding of the 3D scene and can generate high-quality novel views from imperfect conditional information and exhibit robustness to the point cloud conditions.
VI Conclusion
In this paper, we introduce ReconX, a novel sparse-view 3D reconstruction framework that reformulates the inherently ambiguous reconstruction problem as a generation problem. The key to our success is that we unleash the strong prior of video diffusion models to create more plausible observations frames for sparse-view reconstruction. Grounded by the empirical study and theoretical analysis, we propose to incorporate 3D structure guidance into the video diffusion process for better 3D consistent video frames generation. What’s more, we propose a 3D confidence-aware scheme to optimize the final 3DGS from generated frames, which effectively addresses the uncertainty issue. Extensive experiments demonstrate the superiority of our ReconX over the latest state-of-the-art methods in terms of high quality and strong generalizability in unseen data. We believe that ReconX provides a promising research direction to craft intricate 3D worlds from video diffusion models and hope it will inspire more works in the future.
VII Appendix
VII-A Theoretical Proof
Proposition 1. Let $\boldsymbol{\theta}^{*},\psi^{*}=g^{*}$ be the optimal solution of the solely image-based conditional diffusion scheme and $\boldsymbol{\tilde{\theta}}^{*},\tilde{\psi}^{*}=\{{g}^{*},\mathcal{F}^{*}\}$ be the optimal solution of diffusion scheme with native 3D prior. Suppose the divergence $\mathcal{D}$ is convex and the embedding function space $\Psi$ includes all measurable functions, we have $\mathcal{D}(q\left(\boldsymbol{x}\right)\|p_{\boldsymbol{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\boldsymbol{x}))<\mathcal{D}\left(q\left(\boldsymbol{x}\right)\|p_{\boldsymbol{\theta}^{*},\psi^{*}}(\boldsymbol{x})\right)$.
$\mathcal{D}\left(q(\boldsymbol{x})\|p_{\boldsymbol{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\boldsymbol{x})\right)$ $=\mathcal{D}\left(\mathbb{E}_{q(s)}q(\boldsymbol{x}|s)\|\mathbb{E}_{q(s)}p_{\boldsymbol{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\boldsymbol{x}|s)\right)$ (15)
$\leq\mathbb{E}_{q(s)}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\boldsymbol{x}|s)\right)$
$=\mathbb{E}_{q(s)}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{\tilde{\theta}}^{*},{g}^{*},\mathcal{F}^{*}}(\boldsymbol{x}|s)\right),$
where we incorporate an intermediate variable $s$, which represents a specific scene. $q(\boldsymbol{x}|s)$ indicates the conditional distribution of rendering data $\boldsymbol{x}$ given the specific scene $s$. According to the definition of $\boldsymbol{\tilde{\theta}}^{*},{g}^{*},\mathcal{F}^{*}$, we have:
$\mathbb{E}_{q(s)}$ $=\mathbb{E}_{q(s)}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{\tilde{\theta}}^{*},{g}^{*},\mathcal{F}^{*}}(\boldsymbol{x}|s)\right)$ (16)
$=\min_{\boldsymbol{{\theta}},{g},\mathcal{F}}\mathbb{E}_{q(s)}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{{\theta}},{g},\mathcal{F}}(\boldsymbol{x}|s)\right)$
$=\min_{\boldsymbol{{\theta}}}\mathbb{E}_{q(s)}\min_{{g}(s),\mathcal{F}(s)}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{{\theta}},{g}(s),\mathcal{F}(s)}(\boldsymbol{x})\right)$
$=\min_{\boldsymbol{{\theta}}}\mathbb{E}_{q(s)}\min_{{g},E}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{{\theta}},{g},E}(\boldsymbol{x})\right),$
$\mathcal{D}\left(q(\boldsymbol{x})\|p_{\boldsymbol{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\boldsymbol{x})\right)$ $\leq\min_{\boldsymbol{{\theta}}}\mathbb{E}_{q(s)}\min_{{g},E}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{{\theta}},{g},E}(\boldsymbol{x})\right)$ (17)
$<\min_{\boldsymbol{\theta},g,E}\mathcal{D}\left(q(\boldsymbol{x})\|p_{\boldsymbol{\theta},g,E}(\boldsymbol{x})\right)$
$=\min_{\boldsymbol{\theta},g,E(\varnothing)}\mathcal{D}\left(q(\boldsymbol{x})\|p_{\boldsymbol{\theta},g,E(\varnothing)}(\boldsymbol{x})\right)$
$=\min_{\boldsymbol{\theta},\psi}\mathcal{D}\left(q(\boldsymbol{x})\|p_{\boldsymbol{\theta},\psi}(\boldsymbol{x})\right)$
$=\mathcal{D}\left(q\left(\boldsymbol{x}\right)\|p_{\boldsymbol{\theta}^{*},\psi^{*}}(\boldsymbol{x})\right).$
The second inequality holds because given general real-world scene $s$ in any parameter $\boldsymbol{\theta}\in\Theta$, approximating $q(\boldsymbol{x}|s)$ is simpler than $q(\boldsymbol{x})$ by only tuning the encoder $E$ of $p_{\boldsymbol{\theta},g,E}$ [^1] A simple verifiable case is to optimize the parameters of 3DGS by only 2D images (solely image-based conditional learning) or using a SFM initialization from collected images (native 3D conditional learning) before optimization. The latter provides a more constrained and optimal solution space., i.e., $\min_{E}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{{\theta}},{g},E}(\boldsymbol{x})\right)<\min_{E}\mathcal{D}\left(q(\boldsymbol{x})\|p_{\boldsymbol{\theta},g,E}(\boldsymbol{x})\right)$ holds almost everywhere (a.e.), representing $\mathcal{P}_{q(s)}\left\{\min_{E}\mathcal{D}\left(q(\boldsymbol{x}|s)\|p_{\boldsymbol{\theta},g,E}(\boldsymbol{x})\right)<\min_{E}\mathcal{D}\left(q(\boldsymbol{x})\|p_{\boldsymbol{\theta},g,E}(\boldsymbol{x})\right)\right\}$ is equal to 1. Consequently, the proof of Proposition 1 has been done.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China under Grant 62206147.